Given that w denotes either one of the non-real roots of the equation z3 = 1, show that (a) 1 +w+w2 =0, and 3 (b) the other non-real root is w2. Show that the non-real roots of the equation 1 −u u can be expressed in the form Aw and Bw2, where A and B are real numbers, find A and B.
1
Expert's answer
2020-05-20T19:12:15-0400

Given equation,

"z^3=1"

or, "z^3-1=(z-1)(z^2+1+z)=0"

or,"z=1 \ and \ z^2+z+1=0"

Since, z="\omega" is a root of the equation, then, putting the value of z="\omega" in the second equation, we get,

"\omega^2+\omega+1=0" (hence proved)

Now, in the equation,"z^2+z+1=0" , sum of the two roots "=\frac{-b}{a}= -1. \" Since, one root is z="\omega", then the other root will be z= -1-"\omega" . Now from the above relation ("\omega^2+\omega+1=0)" we get that the other root of the equation z=-1-"\omega" can be expressed as z="\omega^2". (proved).

Next, the other equation is "1-u^2=0" . Root of this equation are "u=\frac{+}{-}1". Hence, the roots of this equation can be written as "u=\frac{+}{-}\omega\frac{+}{-}\omega^2." Hence, the values of A and B :

A="\frac{+}{-}1 \ and \ B=\frac{+}{-}1."